OpenStudy (amonoconnor):

Would someone check my work in finding the following anti-derivative? The original expression: "cos((pix)/2)". Any and all help is greatly appreciated! :) I got "(pi/2)(-sin((pix)/2)) + C"

OpenStudy (amonoconnor):

Should I show the work I did?

OpenStudy (shamil98):

yes

OpenStudy (amonoconnor):

Coming right up...

OpenStudy (amonoconnor):

cos[(pix)/2] = (1)cos[(pix)/2] = (1/((pi/2)(x^(1-1)))cos[(pix)/2] = (1/(pi/2))(-sin[(pix)/2] = ((1)(2/pi))-sin[(pix)/2] = (2/pi)-sin[(pix)/2] + C

OpenStudy (shamil98):

\[\int\limits_{}^{} \cos(\frac{ \pi }{ 2}) x~ dx\]

OpenStudy (amonoconnor):

Umm, x(sin(pi/2)) - (sin(pi/2))(1) (?)

OpenStudy (amonoconnor):

Meh, lol, all I wanted to know was if I was right... :)

OpenStudy (shamil98):

Just go with the answer, I can't really confirm if it is correct or not..

OpenStudy (amonoconnor):

Okay... Good deal;)